OpenStudy (anonymous):

During a 4th of July weekend, 32 vehicles became trapped on the Sunshine Skyway Bridge while it was being repaved. A recent city ordinance decreed that only cars with 4 wheels & trucks w/ 6 wheels could be on the bridge at any given time. If there were 148 tires that needed to be replaced to due to damage, how many cars & trucks were involved in the incident?

OpenStudy (anonymous):

Cars + trucks=32 4 cars+6 trucks=148 cars=32-t 4(32-t)+6 trucks=148 128-4 trucks+6 trucks=148 2 trucks=20 trucks=10 cars=22 ??